Circular No. 7755

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

LEONID METEORS 2001 A strong showing of Leonid meteors on Nov. 18 has been reported

by observers in North America and in eastern Asia and Australia. Visual observations by D. W. E. Green (Lexington, MA) and A. Hale (Cloudcroft, NM, despite clouds) indicate a peak around Nov. 18.43 UT with perhaps 5–15 Leonids per minute; there is some indication that this peak extended to ∼ Nov. 18.46. This first peak occurred close to or about half an hour later than the time predicted by Lyytinen et al. (2001, WGN 29, 110) and about half an hour to an hour later than that predicted by McNaught and Asher (2001, WGN 29, 156), based on ejection from comet 55P/Tempel-Tuttle seven revolutions ago. Rates continued at an elevated level of a couple to several Leonids per minute during Nov. 18.5–18.7, according to S. J. O’Meara (Mauna Kea, HI), D. O’Meara (Volcano, HI), K. Mameta (Kobe, Japan), and S. Nakano (Sumoto, Japan). Mameta reports that a peak rate occurred of 55 Leonids per minute occurred over 10 minutes at Nov. 18.764, and Nakano reports that 4–6 Leonids per second were visible around maximum; by Nov. 18.81, Nakano notes that the rate dropped to ∼ 5–10 Leonids per minute. This second peak occurred close to the times predicted by Lyytinen et al. and by McNaught and Asher, based on ejection from 55P four revolutions ago.

SUPERNOVA 2001ga IN ANONYMOUS GALAXY B. Barris, University of Hawaii; and G. Miknaitis, University of Wash-

ington, on behalf of the High-Z Supernova Search Team (cf. IAUC 6160, 6646, 7312, 7516, 7745), report the discovery of a supernova (I = 19.5) on I-band images taken with the Canada-France-Hawaii Telescope (+ 12K Mosaic in queue mode) on Nov. 12; the supernova was identified by subtracting images taken on Sept. 11 from those taken on Nov. 12. The new object is located at α = 2^{h}27^{m}51^{s}.76, δ = +0^{o}37′20′′.9 (equinox 2000.0), which is 1′′.8 east and 2′′.0 north of the apparent host galaxy. Additional I magnitudes: Nov. 17, 19.7 (P. Garnavich and J. Quinn, 1.8-m Vatican Advanced Technology Telescope); Nov. 18, 19.5 (P. Challis and E. Martin, Subaru telescope + SUPRIME). Spectral observations made with the Keck II telescope (+ ESI) on Nov. 18 by A. V. Filippenko and R. Chornock, University of California at Berkeley, indicate that this is a type-Ia supernova at z = 0.13, ≈ 3 days after maximum brightness.

2001 November 19  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2001 CBAT  Daniel W. E. Green